What is the sum of all positive integer values of n such (n+20)/n is an integer ?
Answer by jim_thompson5910(35256)

Hint: Since (n+20)/n = n/n + 20/n = 1 + 20/n, this means that (n+20)/n is only an integer if n is a factor of 20.